OpenStudy (anonymous):

The parabola with focus (0, 5) and directrix y = -5.

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

well actually I just need help with figuring out this equation

OpenStudy (anonymous):

since the focus is at (0,5) and the directrix is at \(y=-5\) you know this parabola opens up and has vertex at the origin (0,0)

OpenStudy (anonymous):

therefore the equation looks like \(4py=x^2\) and since \(p=5\) you get \(20y=x^2\)

OpenStudy (anonymous):

@aussy123 do you have any question?

OpenStudy (anonymous):

where did you get the 4

OpenStudy (anonymous):

From parabola formula ( x -h)² = 4p ( y - k)

OpenStudy (anonymous):

That's the parabola rule to memorize!

OpenStudy (anonymous):

oh i was givin the distance formula but thanks ill write this formula down

OpenStudy (anonymous):

i got it directly from the old school way to write a parabola with vertex \((h,k)\) it is either \((y-k)^2=4p(x-h)\) or \((x-h)^2=4p(y-k)\) depending on how it opens, to the right/left or up/down

OpenStudy (anonymous):

in your case it opens up, and the vertex is at (0,0) so i wrote \(4py=x^2\)

OpenStudy (anonymous):

oh ok that nakes alot of sense thx

OpenStudy (anonymous):

then since \(p=5\) is the focus, we know it is \(4\times 5y=x^2\)